OpenStudy (keke):

simplify; x^(2)-5x over 5x-25

OpenStudy (anonymous):

\[\frac{x^2-5 x}{5 x-25}=\frac{((x-5) x)}{(5 (x-5))}=\frac{x}{5} \]

jhonyy9 (jhonyy9):

(x*2-5x)/(5x-25) =(x(x-5))/(5(x-5))=x/5